# Fast Online Adaptive Neural MPC via Meta-Learning

### v4

@@ 2 Problem Statement
Here, $\ell(x_{k},u_{k})$ is the stage cost, $m(x_{N})$ is the terminal cost, and $\phi(\cdot)$ is a general numerical integrator that discretizes the continuous-time dynamics $f$ over time steps $\delta t=t_{f}/N$. A common example of a numerical integrator is the fourth-order Runge–Kutta method. $N$ denotes the number of discretized steps in the prediction horizon. The constraint $g(x_{k},u_{k})$ models general nonlinear state and input constraints. The optimization is initialized at the current system state $x(t)$ and produces an optimal control sequence $\mathbf{u}_{0:N-1}^{*}$. Only the first control input $u_{0}$ is applied to the system at time $t$, following the receding horizon principle.
@@ 3.1 Robotic System Setup
$m(x_{N})=\left\|x_{N}-x_{\text{ref},k}\right\|_{Q}^{2},$ (5)
@@
where $u_{k}^{(i)}$ denotes the $i$-th control input at time step $k$, and $u_{\min}^{(i)}$, $u_{\max}^{(i)}$ define the lower and upper bounds for each input dimension $i$.
@@ 4.1 Van der Pol Oscillator
[…]
@@ 4.2 Cart-Pole System
[…]
@@
Prediction horizon steps $N$ 20
@@
Figure 6: The comparison results of stabilizing the 2D quadrotor. The shaded regions represent standard deviations over 20 trials. (a) Averaged absolute errors of $x$ over time. (b) Averaged absolute errors of $z$ over time. (c) Averaged absolute Euclidean errors of $x-z$ over time.
@@
$\mathbf{x}_{\text{ref}}(t)=\begin{bmatrix}x_{\text{ref}}(t)\\ \dot{x}_{\text{ref}}(t)\\ z_{\text{ref}}(t)\\ \dot{z}_{\text{ref}}(t)\\ \theta_{\text{ref}}(t)\\ \dot{\theta}_{\text{ref}}(t)\end{bmatrix}=\begin{bmatrix}0.5\cos\left(\frac{2\pi}{15}t\right)\\ 0\\ 1+0.5\sin\left(\frac{2\pi}{15}t\right)\\ 0\\ 0\\ 0\end{bmatrix}.$ (12)
@@
Figure 7: The comparison results of tracking circular reference. The shaded regions represent standard deviations over 20 trials. (a) Averaged absolute errors of $x$ over time. (b) Averaged absolute errors of $z$ over time. (c) Averaged absolute Euclidean errors of $x-z$ over time.

### v5

@@ 2 Problem Statement
Here, $\ell(x_{k},u_{k})$ is the stage cost, $m(x_{N})$ is the terminal cost, and $\phi(\cdot)$ is a general numerical integrator that discretizes the continuous-time dynamics $f$ over time steps $\delta t=t_{f}/N$. A common example of a numerical integrator is the fourth-order Runge–Kutta method. NN denotes the number of discretized steps in the prediction horizon. The constraint $g(x_{k},u_{k})$ models general nonlinear state and input constraints. The optimization is initialized at the current system state $x(t)$ and produces an optimal control sequence $\mathbf{u}_{0:N-1}^{*}$. Only the first control input $u_{0}$ is applied to the system at time $t$, following the receding horizon principle.
@@ 3.1 Robotic System Setup
$m(x_{N})=\left\|x_{N}-x_{\text{ref},N}\right\|_{Q}^{2},$ (5)
@@
where $u_{k}^{(i)}$ denotes the $i$-th control input at time step $k$, and $u_{\min}^{(i)}$, $u_{\max}^{(i)}$ define the lower and upper bounds for each input dimension ii.
@@ 4.1 Van der Pol Oscillator
To best simulate the real-time controller scenario, we use the models from three MPC controllers to predict the trajectory $x_{1},x_{2}]$ over a 1-second prediction horizon. The residual dynamics models (MLP and MetaMLP) are fine-tuned online using observed error from the previous 1-second window, and the observation considers the measurement Gaussian noise $\mathcal{N}(0,\,0.025^{2})$. Given the system operates at 50 Hz, each fine-tuning step uses a batch of 50 samples. Consequently, the MetaMLP is updated with $K=50$ samples per adaptation. Other architecture and hyperparameter settings can be found in Table I. In the meta-training phase, we collected tasks with varying $\mu$ values, sampled in increments of 0.1 from 0 to 1. For each task, we simulated the oscillator dynamics starting from a random initial condition $x_{0}\in[-2,2]^{2}$ over a duration of 10 seconds. The initial model parameters $\theta_{0}$ were trained using the meta-learning approach described in Section 3.3, with the hyperparameters listed in Table I(b).
@@ 4.2 Cart-Pole System
$\ddot{\theta}$ $=\frac{g\sin\theta+\cos\theta\left(\frac{-F-m_{p}l\dot{\theta}^{2}\sin\theta}{m_{c}+m_{p}}\right)}{l\left(\frac{4}{3}-\frac{m_{p}\cos^{2}\theta}{m_{c}+m_{p}}\right)},$
@@
Prediction horizon steps NN 20
@@
Figure 6: The comparison results of stabilizing the 2D quadrotor. The shaded regions represent standard deviations over 20 trials. (a) Averaged absolute errors of $x$ over time. (b) Averaged absolute errors of $z$ over time. (c) Averaged absolute Euclidean errors of xzx-z over time.
@@
$\mathbf{x}_{\text{ref}}(t)=\begin{bmatrix}x_{\text{ref}}(t)\\ \dot{x}_{\text{ref}}(t)\\ z_{\text{ref}}(t)\\ \dot{z}_{\text{ref}}(t)\\ \theta_{\text{ref}}(t)\\ \dot{\theta}_{\text{ref}}(t)\end{bmatrix}=\begin{bmatrix}0.5\cos\!\left(\frac{2\pi}{15}\,t\right)\\ -\frac{\pi}{15}\,\sin\!\left(\frac{2\pi}{15}\,t\right)\\ 1+0.5\sin\!\left(\frac{2\pi}{15}\,t\right)\\ \frac{\pi}{15}\,\cos\!\left(\frac{2\pi}{15}\,t\right)\\ 0\\ 0\end{bmatrix}$. (12)
@@
Figure 7: The comparison results of tracking circular reference. The shaded regions represent standard deviations over 20 trials. (a) Averaged absolute errors of $x$ over time. (b) Averaged absolute errors of $z$ over time. (c) Averaged absolute Euclidean errors of xzx-z over time.